• This is the second item of the list. It contains another list nested inside it. The inner list is an enumerated list.

1. This is the first item of an enumerated list that is nested within the itemized list.

2. This is the second item of the inner list. L^{A}TEX allows you to nest lists deeper than you really should.

This is the rest of the second item of the outer list. It is no more interesting than any other part of the item.

• This is the third item of the list.

You can even display poetry.

There is an environment for verse Whose features some poets will curse.

For instead of making Them do all line breaking, It allows them to put too many words on a line when

they’d rather be forced to be terse.

Mathematical formulas may also be displayed. A displayed formula is one-line long; multiline formulas require special formatting instructions.

(Γ, ψ′) = x′′ + y^{2} + z_{i}^{n}

Don’t start a paragraph with a displayed equation, nor make one a paragraph by itself.

An Example Document

Leslie Lamport

January 21, 1994

This is an example input file. Comparing it with the output it generates can show you how to produce a simple document of your own.

1 Ordinary Text

The ends of words and sentences are marked by spaces. It doesn’t matter how many spaces you type; one is as good as 100. The end of a line counts as a space.

One or more blank lines denote the end of a paragraph. Since any number of consecutive spaces are treated like a

single one, the formatting of the input file makes no difference to L^{A}TEX, but it makes a difference to you. When you use L^{A}TEX, making your input file as easy to read as possible will be a great help as you write your document and when you change it. This sample file shows how you can add comments to your own input file.

1